OpenStudy (anonymous):

given the formula, solve for the indicated variable A=P(1+nr) for r

OpenStudy (paxpolaris):

star by dividing both sides by P:\[\frac AP=1+n \cdot r\] then subtract the 1: \[\implies \frac AP -1 = n \cdot r\]\[\implies {A -P \over P} = n \cdot r\] Lastly divide both sides by n \[\implies \Large r={A -P \over P \cdot n}\]